OpenStudy (anonymous):

I have a homework question! I have no idea what to do.. Please help! Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x2 − 4x + y2 + 8y = −4

OpenStudy (anonymous):

well do u knw wats the general equation of circle .. u have to compare this with general form

OpenStudy (anonymous):

x2-4x+4 +y2+8y+16-4-16=-4

OpenStudy (anonymous):

just added and subtracted 4 and 16

OpenStudy (anonymous):

(x2-4x+4 )+(y2+8y+16)-4-16=-4

OpenStudy (anonymous):

(x-2)^2+(y+4)^2-20=-4

OpenStudy (anonymous):

(x-2)^2+(y+4)^2=16 (x-2)^2+(y+4)^2=4^2

OpenStudy (anonymous):

standard for of circle equation is (x-h)^2+(y-k)=r^2 where (h,k) is circle`s center and r is radius so compare it with (x-2)^2+(y+4)^2=4^2

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

u are welcoem .. so tell me wats ur center and wats ur radius ????